### v2

@@ 5.1.2 Minimal coupling $\xi_{\chi}=0$
In this case, the derivative potential $V_{\phi}$ and the Hubble parameter $H^{2}$ contribute so that the effective mass $m_{\mathrm{end}}$ is negative. Thus, once the modes cross the horizon and for masses $m_{\chi}\ll H_{\mathrm{end}}$, this may lead to $\omega_{k}^{2}(t)<0$, resulting in tachyonic instabilities, as mentioned in the discussion following equation (41) [41]. These instabilities primarily affect the IR modes of the produced particle spectrum, as the modes outside the horizon during inflation experience tachyonic growth. This growth of the modes $X_{\mathbf{k}}(t)$ outside the horizon is the main driver of the enhanced particle production for minimally coupled scalars compared to conformally coupled scalars. The spectrum of the produced particles is also affected, and its behavior depends on the mass $m_{\chi}$. To obtain the gravitational spectra, we use the same technique to solve the equation numerically: we choose the $k$-mode inside the horizon, approximately 5 e-folds before it crosses the horizon and show the numerical result for $\beta_{k}(t)$ at those times when $\omega_{k}^{2}>0$. In the figure 8, we show the spectrum of created particles for minimally coupled masses. We observe that the spectrum is nearly scale-invariant for IR modes and for masses $m_{\chi}\lesssim H_{\mathrm{end}}$ before reaching the peak at $k_{\mathrm{peak}}$. However, as the mass increases, the spectrum tends to be blue-tilted. In the UV region, all modes decay either as a power law or exponentially.
Figure 9: Comoving number density of particles as a function of cosmic time. A similar behavior can be observed in Figure 6: the number of created particles saturates at the end of inflation and stabilizes, becoming constant at late times
A clear difference exists between the cases $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$. In the frequency $\omega_{k}^{2}(t)$, the Ricci scalar does not appear for $\xi_{\chi}=0$, indicating a significant difference in the behavior of $\omega_{k}$ for different coupling cases. This is because the Ricci scalar exhibits an oscillatory behavior after inflation due to the oscillations of the fields $\phi$ and $h$. In other words, for the case $\xi_{\chi}=1/6$, the Ricci scalar plays an important role in describing the evolution of the modes $X_{k}(t)$, which translates into a different spectrum of created particles. From figure 9 we can see that the particle production at minimal coupling is much larger than the conformally-coupled case. This is mainly due to the behavior of the spectator field $X_{k}(t)$ during inflation. As explained above, the tachyonic instabilities in the case of $\xi_{\chi}=0$ cause the modes $X_{k}(t)$ to grow exponentially. This growth of the super-horizon modes is primarily responsible for the enhanced gravitational particle production for minimally coupled scalars.
@@ 6 Discussion and remarks
In this study, we investigated the gravitational and non-perturbative production of free scalar particles, $\chi$, in Higgs-$R^{2}$ inflation. We started with a free spectator field, $\chi$, interacting only through the gravitational term $\xi_{\chi}R$ in the Jordan frame. By applying a conformal transformation to transition into the Einstein frame, direct couplings between the $\chi$ field and the scalaron $\phi$ naturally emerged. Within the context of inflationary dynamics in the Einstein frame, we derived the equation of motion for the $\chi$ field, accounting for its non-minimal coupling to the scalar curvature $R$. We numerically solved this system of equations, imposing Bunch-Davies conditions on the $\chi$ field five e-folds before the modes crossed the horizon, to determine the mode functions of $\chi$. From these solutions, we extracted the Bogolyubov coefficient $\beta_{k}(t)$ to compute the comoving particle density and visualize the resulting spectrum. We explored two distinct scenarios for the non-minimal coupling parameter $\xi_{\chi}$, considering different mass ranges. For $\xi_{\chi}=1/6$, we observed efficient particle production at the end of inflation, driven by the term $e^{-\alpha\phi}m_{\chi}^{2}$, which becomes dominant at this stage. The resulting power spectrum exhibited a blue-tilted behavior, indicating that particle production was dominated by modes near the Hubble radius $(aH)^{-1}$ at the end of inflation, while shorter-wavelength modes were suppressed. As the mass increased in the $\xi_{\chi}=1/6$ case, particle production tended to decrease due to the increase in the effective mass. In contrast, for $\xi_{\chi}=0$, the spectrum remained nearly scale-invariant for masses $m_{\chi}\lesssim H_{\mathrm{end}}$, with a blue tilt emerging for larger masses $m_{\chi}\gtrsim H_{\mathrm{end}}$. For light masses and $\xi_{\chi}=0$, we found that the comoving number of particles created was significantly higher compared to the case of conformally coupled particles, primarily due to the tachyonic growth of the modes $X_{k}(t)$.
@@
We did not attempt a detailed exploration of the parameter space of the Higgs-$R^{2}$ inflation model or the parameters of the scalar field $\chi$. Instead, we focused on the Higgs-$R^{2}$ inflation model, whose observables are consistent with Planck data, and considered the possibility of non-perturbative particle production through the direct couplings that arise between the $\chi$ field and the inflationary sector when changing the conformal frame. We found that inflation driven by the Higgs field and the scalaron can produce scalar particles in the cases $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$ for a specific range of masses. This field could potentially contribute to the abundance of dark matter, though its present-day abundance would need to be calculated. There remain several directions for future research, including the production of spin-$1/2$ and spin-$1$ dark matter, the production of gauge bosons $W^{\pm}_{\mu}$, $Z_{\mu}$, and the implications of particle production on primordial non-Gaussianities in the CMB. It would also be interesting to study the isocurvature constraints for light scalar field masses in this model. We leave these intriguing topics for future investigation.
@@ Acknowledgements
We are very grateful to Dr. Marcos Garcia for his insightful comments and input, which made it possible to write the code for this article. We would also like to acknowledge Dr. Andrew Long for providing an example of numerical code. The work of F.P. has been sponsored by CONACHyT-Mexico through a doctoral scholarship. This work was also partially supported by CONACHyT- Mexico through the support CBF2023-2024-193

### v3

@@ 5.1.2 Minimal coupling $\xi_{\chi}=0$
In this case, the derivative potential $V_{\phi}$ and the Hubble parameter $H^{2}$ contribute so that the effective mass $m_{\mathrm{end}}$ is negative. During inflation where $\phi\simeq 5.7M_{p}$, $\dot{H}\to 0$ and the mass term $e^{-\alpha\phi}m_{\chi}^{2}\ll H$, so it is negligible. In these cases, the dominant term of (50) is $H^{2}$ if the modes are super-Horizon $k\ll aH$. Thus, once the modes cross the horizon and for masses $m_{\chi}\ll H_{\mathrm{end}}$, this may lead to $\omega_{k}^{2}(t)<0$, resulting in tachyonic instabilities, as mentioned in the discussion following equation (36) [46]. These instabilities primarily affect the IR modes of the produced particle spectrum, as the modes outside the horizon during inflation experience tachyonic growth. This growth of the modes $X_{\mathbf{k}}(t)$ outside the horizon is the main driver of the enhanced particle production for minimally coupled scalars compared to conformally coupled scalars.
Figure 9: The Bogolyubov coefficient $|\beta_{k}|^{2}$ as a function of cosmic time $t$ for a spectator mass $m_{\chi}=3H_{\mathrm{end}}$ and minimal coupling $\xi_{\chi}=0$.
The spectrum of the produced particles is also affected, and its behavior depends on the mass $m_{\chi}$. To obtain the gravitational spectra, we use the same technique to solve the equation numerically: we choose the $k$-mode inside the horizon, approximately 5 e-folds before it crosses the horizon and show the numerical result for $\beta_{k}(t)$ at those times when $\omega_{k}^{2}>0$. In Fig. 8, we show the spectrum of created particles for minimally coupled masses. We observe that the spectrum is nearly scale-invariant for IR modes and for masses $m_{\chi}\lesssim H_{\mathrm{end}}$ before reaching the peak at $k_{\mathrm{peak}}$. However, as the mass increases, the spectrum tends to be blue-tilted. In the UV region, all modes decay either as a power law or exponentially [49]. As the authors mention, the comoving particle number density (45) can have an IR divergence for $m_{\chi}\ll H_{\mathrm{end}}$ that is governed by a cutoff scale $k_{0}$. For $\xi_{\chi}>0$, the integral (45) converges and becomes IR insensitive approximately when $\xi_{\chi}\sim 1/6$. A clear difference exists between the cases $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$. In the frequency $\omega_{k}^{2}(t)$, the Ricci scalar does not appear for $\xi_{\chi}=0$, indicating a significant difference in the behavior of $\omega_{k}$ for different coupling cases. This is because the Ricci scalar exhibits an oscillatory behavior after inflation due to the oscillations of the fields $\phi$ and $h$. In other words, for the case $\xi_{\chi}=1/6$, the Ricci scalar plays an important role in describing the evolution of the modes $X_{k}(t)$, which translates into a different spectrum of created particles. In the Fig. 9 we present the behavior of $|\beta_{k}(t)|^{2}$ as a function of time for a mass $m_{\chi}=3H_{\mathrm{end}}$ and $N_{k}=60$ for the minimum coupling case $\xi_{\chi}=0$. We see the same behavior as in the case $\xi_{\chi}=1/6$, a resonant peak near the end of inflation, subsequent oscillations until reaching a stabilized and constant value at $t\gg t_{\mathrm{end}}$. The difference lies in the order of magnitude of $|\beta_{k}(t)|^{2}$, and hence of $n_{k}$. We can see that the particle production at minimal coupling is much larger than the conformally-coupled case. This is mainly due to the behavior of the spectator field $X_{k}(t)$ during inflation.
As explained above, the tachyonic instabilities in the case of $\xi_{\chi}=0$ cause the modes $X_{k}(t)$ to grow exponentially. This growth of the super-horizon modes is primarily responsible for the enhanced gravitational particle production for minimally coupled scalars [49, 5].
6 Relic abundance
If gravitationally produced particles are sufficiently long-lived, they could survive throughout the entire thermal history of the universe up to the present-day $t=t_{0}$, thereby becoming viable dark matter candidates. These particles might account for only a fraction of the observed dark matter, or they could potentially explain its entirety, as quantified by the parameter $\Omega_{\mathrm{DM}}$. For this reason, it is worthwhile to compute both the total number density $n_{\chi}$ and the present-day abundance $\Omega_{\chi}h^{2}$. In this section, we compute the present-day relic abundance of $\chi$ particles produced gravitationally in the Higgs-$R^{2}$ model, summarize our numerical results, and impose the corresponding relic abundance constraints. The total number of gravitationally produced particles is characterized by two key parameters, depending on the application: either through the comoving particle number density, defined in Eq. (45), or via the present-day density parameter $\Omega_{\chi}$. The present-day particle relic abundance is given by
$\Omega_{\chi}h^{2}=\dfrac{\rho_{\chi}(t_{0})}{\rho_{c}}h^{2}\,.$ (51)
Figure 10: Total comoving number density $a^{3}n_{\chi}$ in units of $k_{\mathrm{end}}$ as a function of the spectator mass $m_{\chi}/H_{\text{end}}$. The left panel shows the result for the conformal coupling case $\xi_{\chi}=1/6$, while the right panel shows the result for minimal coupling $\xi_{\chi}=0$.
Here, $\rho_{\chi}$ is the total energy density of the produced particles, and $\rho_{c}=3M_{p}^{2}H_{0}^{2}$ is the critical density of the universe today. Once the asymptotic value of $n_{k}(t)$ is known, we can compute $\Omega_{\chi}h^{2}$ at the present-day. The renormalized total energy density $\rho_{\chi}$ is defined as [5, 9]:
$\rho_{\chi}(t)a^{4}(t)=\int\dfrac{\mathrm{d}^{3}k}{(2\pi)^{3}}\,\omega_{k}(t)\,|\beta_{k}(t)|^{2}=\int_{0}^{\infty}\dfrac{\mathrm{d}k}{k}\,\omega_{k}(t)n_{k}(t)\,.$ (52)
For non-relativistic modes $k\ll a\,m_{\chi}$, which are expected to carry most of the energy at late times, we have:
$\rho_{\chi}(t)a^{4}(t)\simeq\dfrac{m_{\chi}n_{\chi}}{\rho_{c}}\,.$ (53)
That is, the present-day abundance at $t=t_{0}$ can be computed from:
$\Omega_{\chi}h^{2}\simeq\dfrac{m_{\chi}n_{\chi}(t_{0})}{\rho_{c}}h^{2}\,.$ (54)
It is important to emphasize that the mass term $m_{\chi}$ appearing in the frequency (34) is accompanied by an exponential factor $e^{-\alpha\phi}$, and thus it must be evaluated once $\phi$ has stabilized after reheating. To perform this calculation, one needs to know the reheating mechanism, the reheating temperature $T_{\mathrm{RH}}$, and its duration. However, our particle production calculation is independent of the reheating process, since we assume that the scalar field is a spectator. Reheating should take into account the production of gauge bosons $W$ and $Z$ due to Higgs decay and possible resonances.
First, we show the total comoving number density $a^{3}n_{\chi}$ as a function of $m_{\chi}$ for both $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$ in Fig. 10. In both cases, we observe a decrease in the number of produced particles as the mass increases, as expected. The superheavy case $m_{\chi}\gg H_{\text{end}}$ is particularly interesting for minimal coupling. In the right panel of Fig. 10, we observe a significant enhancement in the comoving number density $a^{3}n_{\chi}$ over the explored mass range, reaching a maximum of order $10^{5}$ for $m_{\chi}/H_{\text{end}}\sim 2.5$, which corresponds to $m_{\chi}\simeq 1.5\times 10^{13}\,\text{GeV}$. Although this mass is not strictly in the $m_{\chi}\gg H_{\text{end}}$ regime, as the mass increases further, $a^{3}n_{\chi}$ is expected to decay. However, for a more detailed study of the $\xi_{\chi}=0$ case, it is necessary to extend the mass range to the superheavy case, which is computationally more demanding, and requires a much finer numerical analysis. This is left for future work. The conformally coupled case $\xi_{\chi}=1/6$ follows the expected behavior: it exhibits a peak around $m_{\chi}\sim H_{\text{end}}$, and for heavier spectator masses $m_{\chi}\gtrsim H_{\text{end}}$, gravitational production becomes suppressed.
To calculate the $\Omega_{\chi}$ abundance, it is necessary to understand the mechanism of the reheating in the model. As mentioned, our work is independent of this process. There are several attempts to explain ths epoch in the Higgs-$R^{2}$ model. For example, in [37], the non-perturbative production of $W$ bosons and Higgs boson self-production after inflation is analyzed. This allows for the computation of relevant observables such as the reheating temperature $T_{\mathrm{rh}}$. The authors conclude that the dominant post-inflationary dynamics of the model is driven by tachyonic instabilities in the effective masses of the gauge bosons, and to a lesser extent, by the tachyonic mass of the Higgs boson itself. This tachyonic regime leads to a cosmologically instantaneous preheating, which allows the post-inflationary evolution of the model to be fully determined. In light of these results, we can provide an illustrative analytical estimate for the dark matter abundance $\Omega_{\chi}h^{2}$ of the spectator field $\chi$, based on the comoving number density $n_{\chi}$ and the reheating temperature $T_{\text{RH}}$. We evaluate the total comoving number density of produced particles at late times, taken as the final time $t_{\text{final}}$ of our numerical simulations, where the factor $a^{3}n_{\chi}$ has reached a constant value. Using the asymptotic value of $n_{\chi}$, we can calculate the relic abundance $\Omega_{\chi}h^{2}$ using the amount of expansion from the end of inflation to today. To do this, we must rescale the scale factor from the end of inflation to the present-day, taking into account the reheating epoch. Taking into account that after inflation, the production of gauge bosons begins due to the decay of the Higgs field [37], the expression (54) can be rewritten as
$\Omega_{\chi}h^{2}\simeq\dfrac{m_{\chi}n_{\chi}(t_{0})}{\rho_{c}}=\dfrac{m_{\chi}}{3M_{p}^{2}H_{0}^{2}}\dfrac{a^{3}n_{\chi}}{k_{\text{end}}^{3}}\dfrac{k_{\text{end}}^{3}}{a_{0}^{3}}\,,$ (55)
where $n_{\chi}(t_{0})=a^{3}n_{\chi}/k_{\text{end}}^{3}=H_{\text{end}}^{3}[a^{3}n_{\chi}/k_{\text{end}}^{3}](a_{\text{end}}/a_{0})^{3}$ is the comoving number density at the present-day. The factor $[a^{3}n_{\chi}/k_{\text{end}}^{3}]$ comes from calculating (45) numerically, see Fig. 10. The ratio of the scale factors $(a_{\text{end}}/a_{0})^{3}$ corresponds to the amount of expansion from the end of inflation to the present-day. Taking into account that after reheating the entropy density $s$ is conserved, the amount of expansion from the end of inflation to the present time can be rewritten as [51, 11]
Figure 11: Dimensionless relic abundance (54) of $\chi$ particles for the case $\xi_{\chi}=1/6$ as a function of the mass, assuming a reheating temperature $T_{\text{RH}}\sim 10^{9}\,\text{GeV}$. The horizontal line indicates the present-day value of the dark matter abundance $\Omega_{\text{CDM}}h^{2}\sim 0.12$.
$\left(\dfrac{a_{\text{end}}}{a_{0}}\right)^{3}=\dfrac{\pi^{2}g_{*}T_{0}^{3}}{90M_{p}^{2}H_{\text{end}}^{2}}T_{\text{RH}}\,.$ (56)
Therefore, the abundance (54) of the created particles rescaled to the present-day becomes
$\Omega_{\chi}h^{2}=\left(\dfrac{\pi^{2}g_{*}T_{0}^{3}}{270M_{p}^{4}H_{0}^{2}}\right)\left(\dfrac{m_{\chi}}{H_{\text{end}}}\right)\left(\dfrac{a^{3}n_{\chi}}{k_{\text{end}}^{3}}\right)H_{\text{end}}^{2}\,T_{\text{RH}}\,,$ (57)
where $g_{*}\approx 3.91$, $T_{0}=0.234\times 10^{-12}\,\text{GeV}$ and $H_{0}=2.133\times 10^{-42}\,\text{GeV}$ are the numerical values in GeV of the temperature and the Hubble constant at the present-day. Thus, the numerical factor of $\Omega_{\chi}h^{2}$ is
$\dfrac{\pi^{2}g_{*}T_{0}^{3}}{270M_{p}^{4}H_{0}^{2}}\approx 0.1154\times 10^{-28}\,\text{GeV}^{-3}$ (58)
The expression (57) allows us to calculate the relic abundance from the numerical calculation of the comoving number density $a^{3}n_{\chi}$, and as a function of the mass $m_{\chi}$ and the reheating temperature $T_{\text{RH}}$. For example, by fixing $T_{\mathrm{RH}}$, one can estimate the dependence of $\Omega_{\chi}$ on the mass $m_{\chi}$; alternatively, by fixing the ratio $m_{\chi}/H_{\text{end}}$, the abundance becomes proportional to $a^{3}n_{\chi}/k_{\mathrm{end}}^{3}$, isolating the effect of particle production. Since the reheating temperature is a fixed prediction of this model, we focus on exploring the dependence of $\Omega_{\chi}$ on the mass $m_{\chi}$. In Fig. 11, we present the relic abundance of $\chi$ particles at the present-day as a function of their mass for $\xi_{\chi}=1/6$, using Eq. (57) and $T_{\text{RH}}\sim 10^{9}$ GeV. From the figure we find that the maximum allowed mass for the spectator field to account for the observed value $\Omega_{\text{dm}}h^{2}$ without overproduction is $m_{\chi}\simeq 10^{12}\,\text{GeV}$, consistent with other results in the literature for the conformally coupled case [8, 13], although the exact upper bound may vary depending on the specific production mechanism considered. We also observe that for light particles $m_{\chi}\ll H_{\text{end}}$, the abundance decreases. This is because in this regime the gravitational production is inefficient due to the conformal coupling, which suppresses the amplification of modes during the expansion of the Universe. Although these masses fail to explain all of the observed dark matter, they may contribute a subdominant fraction. Therefore, the allowed masses are in the range $m_{\chi}\lesssim 10^{12}\,\text{GeV}$ for the conformal case, considering only non-perturbative dark matter production and heavy particles $m_{\chi}\gtrsim H_{\text{end}}$.
@@ 7 Discussion and remarks
In this paper, we have investigated the gravitational and non-perturbative production of free scalar particles, $\chi$, in Higgs-$R^{2}$ inflation. We started with a free spectator field, $\chi$, interacting only through the gravitational term $\xi_{\chi}R$ in the Jordan frame. By applying a conformal transformation to transition into the Einstein frame, direct couplings between the $\chi$ field and the scalaron $\phi$ naturally emerged. Within the context of inflationary dynamics in the Einstein frame, we derived the equation of motion for the $\chi$ field, accounting for its non-minimal coupling to the scalar curvature $R$. We numerically solved this system of equations, imposing Bunch-Davies conditions on the $\chi$ field five e-folds before the modes crossed the horizon, to determine the mode functions of $\chi$. From these solutions, we extracted the Bogolyubov coefficient $\beta_{k}(t)$ to compute the comoving particle density and visualize the resulting spectrum. We explored two distinct scenarios for the non-minimal coupling parameter $\xi_{\chi}$, considering different mass ranges. For $\xi_{\chi}=1/6$, we observed efficient particle production at the end of inflation, driven by the term $e^{-\alpha\phi}m_{\chi}^{2}$, which becomes dominant at this stage. The resulting power spectrum exhibited a blue-tilted behavior, indicating that particle production was dominated by modes near the Hubble radius $(aH)^{-1}$ at the end of inflation, while shorter-wavelength modes were suppressed. As the mass increased in the $\xi_{\chi}=1/6$ case, particle production tended to decrease due to the increase in the effective mass. In contrast, for $\xi_{\chi}=0$, the spectrum remained nearly scale-invariant for masses $m_{\chi}\lesssim H_{\mathrm{end}}$, with a blue tilt emerging for larger masses $m_{\chi}\gtrsim H_{\mathrm{end}}$. For light masses and $\xi_{\chi}=0$, we found that the comoving number of particles created was significantly higher compared to the case of conformally coupled particles, primarily due to the tachyonic growth of the modes $X_{k}(t)$.
While we did not conduct a comprehensive scan of the parameter space for Higgs-$R^{2}$ inflation or the $\chi$ field, we focused on a benchmark scenario consistent with current CMB observations. We found that non-perturbative production of spectator scalars with conformal coupling can yield viable dark matter candidates for masses up to $m_{\chi}\sim 10^{12}\,\mathrm{GeV}$. Lighter masses fail to account for the full relic abundance but may still contribute a subdominant component. The superheavy regime $m_{\chi}\gg H_{\text{end}}$ is particularly intriguing, involving a different production mechanism driven by adiabaticity violation, but is left for future work due to its computational demands.
Our results underscore the importance of gravitational production mechanisms in extended inflationary scenarios and highlight the viability of free spectator scalars as dark matter candidates, even in the absence of direct couplings to the Standard Model.
@@
There remain several directions for future research, including the production of spin-$1/2$ and spin-$1$ dark matter, the production of gauge bosons $W^{\pm}_{\mu}$, $Z_{\mu}$, and the implications of particle production on primordial non-Gaussianities in the CMB. It would also be interesting to study the isocurvature dark matter constraints for light scalar field masses in this model. We leave these intriguing topics for future investigation.
@@ Acknowledgements
We are very grateful to Dr. Marcos García for his insightful comments and guidance, which were instrumental in the development of the numerical code used in this work. We also thank Dr. Andrew Long for generously providing an example of numerical code that served as a valuable reference during the initial stages of this project. We would like to express our sincere appreciation to the anonymous referee for their constructive suggestions and careful reading of the manuscript, which helped us significantly improve the clarity and quality of this work.
The work of F.P. has been supported by CONACyT-México through the doctoral scholarship 803062. This research was also partially supported by CONACyT-México through the grant CBF2023-2024-193.